One molecule of NaBH4 can in theory reduce four ketone functional groups. If you had 25 mg of NaBH4, how many many mg of benzil could be fully reduced?
Answer #1

Moles of nabh4 = 25*10-3g/ 37.83g/mol =0.66*10-3 moles

This will react  completely with its 4times ketone group (benzil)

Thus moles of benzil = 4*0.66*10-3 = 2.64*10-3 moles

Weight of benzil in these moles ,

Moles = weight /molar mass

=0.00264g/ 210.23g/mol = 0.00001257 g

= 0.01257 mg of benzil